# WonderVerse: Extendable 3D Scene Generation with Video Generative Models

### v2

@@ 2.1 Novel View Synthesis
Novel view synthesis (NVS) is a fundamental vision task to produce a 3D scene from multi-view input images, allowing the rendering of images from novel views. Some traditional methods tackle the problem by formulating 3D scenes as light fields [9, 23], multi-view depth images [15], or blending of a collection of input images [38]. Rather than relying on handcrafted heuristics for constructing 3D scenes, the community later adopts deep learning techniques for NVS [17, 30]. Among these approaches, NeRF [30] has achieved success in improving the image’s quality, resulting in an explosion of follow-up methods [2, barron2021mip, 57, 3, 34, 29] Afterwords, 3D Gaussian Splatting (3DGS) [21] is introduced as a 3D representation that enables fast training and real-time rendering while maintaining high-quality reconstruction results. This has motivated a series of follow-up works [20, 55, 10] that aim to further improve the quality and efficiency of the 3DGS representation. Recently, instead of employing an expensive optimization process, DUSt3R [41] innovatively introduces a pointmap representation, enabling end-to-end 3D reconstruction with fast speed and promising performance. In this work, we demonstrate that our approach is compatible to different 3D reconstruction methods, such as 3DGS and DUSt3R.

### v3

@@ 2.1 Novel View Synthesis
Novel view synthesis (NVS) is a fundamental vision task to produce a 3D scene from multi-view input images, allowing the rendering of images from novel views. Some traditional methods tackle the problem by formulating 3D scenes as light fields [10, 24], multi-view depth images [16], or blending of a collection of input images [40]. Rather than relying on handcrafted heuristics for constructing 3D scenes, the community later adopts deep learning techniques for NVS [18, 31]. Among these approaches, NeRF [31] has achieved success in improving the image’s quality, resulting in an explosion of follow-up methods [2, barron2021mip, 59, 4, 35, 30] Afterwords, 3D Gaussian Splatting (3DGS) [22] is introduced as a 3D representation that enables fast training and real-time rendering while maintaining high-quality reconstruction results. This has motivated a series of follow-up works [21, 57, 11] that aim to further improve the quality and efficiency of the 3DGS representation. Recently, instead of employing an expensive optimization process, DUSt3R [43] innovatively introduces a pointmap representation, enabling end-to-end 3D reconstruction with fast speed and promising performance. In this work, we demonstrate that our approach is compatible to different 3D reconstruction methods, such as 3DGS and DUSt3R.
@@ 5 Conclusion and Limitation
Supplementary Material
6 Additional Comparisons with Existing Works
In this section, we provide additional comparison results with LucidDreamer [9], WonderWorld [55], and Text2Room [20]. As shown in the Fig 89, As visually demonstrated in Figures 8 and 9, our WonderVerse framework can generate extendable 3D scenes, achieving a significant improvement in quality, plausibility, and geometric coherence compared to existing methods.
Baseline approaches that rely on discrete image generation pipelines exhibit noticeable geometric inconsistencies. For instance, LucidDreamer produces fragmented and discontinuous 3D scenes, while Text2Room results in largely distorted 3D scenes, both exhibiting significant geometric inconsistency. WonderWorld, while improved, still exhibits discontinuities between subscenes generated iteratively, which compromises its overall visual quality and realism. On the contrary, WonderVerse produces extendable 3D scenes that are not only visually superior but also maintain geometric integrity, resulting in more believable and immersive 3D environments. This visual comparison underscores the advantage of our approach in creating extendable and coherent 3D scenes.
Figure 8: Qualitative comparison with existing works for extendable 3D scene generation.
Figure 9: Qualitative comparison with existing works for extendable 3D scene generation.
7 Additional Qualitative Results
We show additional qualitative results in Fig 11 13 15 17. We also provide the visualization of their multi-viewed point clouds as shown in Fig 10 12 14 16.
As shown in Figure 10, the reconstructed point clouds from two viewpoints capture the structural elements of an art studio, including walls, windows, and scattered objects. Figure 11 then showcases the rendered scene from multiple perspectives, effectively portraying a vibrant art studio environment populated with canvases and art supplies, consistent with the input text prompt.
Similarly, Figure 12 visualizes the 3D point cloud representing the garden’s layout, capturing the fountain and surrounding floral arrangements. Figure 13 demonstrates the rendered scene, which vividly realizes a colorful flower garden with a central fountain reflecting the sky, faithfully adhering to the descriptive text.
Besides, Figure 14 shows the point cloud reconstruction of the Taj Mahal and its surrounding environment, clearly outlining the iconic structure and pool. Figure 15 showcases the rendered scene, capturing the majestic Taj Mahal with its reflecting pool and cypress trees, demonstrating our framework’s ability to generate complex architectural scenes from textual descriptions.
Beyond generating realistic 3D scenes, our WonderVerse demonstrates the ability to create imaginative scenes that do not exist in the real world, and in diverse styles. As seen in Figure 16, the point cloud reconstruction effectively captures the intricate structure of a Lego space station, including distinct modules, astronauts, and spacecraft elements against a backdrop of stars. Figure 17 then presents the rendered scene, vividly bringing to life a futuristic Lego space station in orbit. This example, rendered in a distinct LEGO style, illustrates our framework’s capability to move beyond photorealistic scene generation. It demonstrates that WonderVerse can interpret textual descriptions to create stylized and fictional 3D environments, expanding its creative potential beyond realistic simulations.
Across these diverse examples, our method consistently generates geometrically plausible and visually compelling 3D scenes that align well with the provided text prompts.
Figure 10: Reconstructed point clouds of our WonderVerse.
Figure 11: Rendered scene of our WonderVerse.
Figure 12: Reconstructed point clouds of our WonderVerse.
Figure 13: Rendered scene of our WonderVerse.
Figure 14: Reconstructed point clouds of our WonderVerse.
Figure 15: Rendered scene of our WonderVerse.
Figure 16: Reconstructed point clouds of our WonderVerse.
Figure 17: Rendered scene of our WonderVerse.
Figure 18: WonderVerse’s interactive generation (a), guided by new text prompts, creates richer 3D scenes compared to non-interactive generation (b), which uses a single initial prompt for all subscenes.
8 Interactive Scene Generation
WonderVerse also enables interactive 3D scene generation. Starting with an initial video based on a text prompt, users can interactively guide scene extension by providing new text prompts to add desired objects. For example, as shown in Figure 18 (a), we interactively extended the scene by adding “A huge fountain” on the right side and “A huge library” on the left side, both successfully incorporated into the 3D scene. Compared to Figure 18 (b), where the entire scene is generated from a single initial prompt, the interactive approach in Figure 18 (a) produces a richer scene aligned with user’s prompts. This example showcases how WonderVerse empowers users to interactively create their own customized and expandable 3D scenes.
9 Details on Efficient 3D Reconstruction with DUSt3R
With DUSt3R [43] as our 3D reconstructor, we capture 25 frames of each video and feed them to DUSt3R and create point cloud with corresponding camera poses. Afterwards, we adopt ICP [3, 37] to register the sub-scene point cloud and finally output a point cloud containing the entire 3D scene.